OpenStudy (anonymous):

Identify the solutions of 3x^2 − 11x − 4 = 0.

OpenStudy (anonymous):

simply solve for x

OpenStudy (anonymous):

3x2 - 11x -4 = 0 3x2 -12x +x -4= 0 now find the solution

OpenStudy (anonymous):

Use the quadratic formula or factor out the problem.

OpenStudy (anonymous):

x=4/3

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@rizwan_uet

OpenStudy (anonymous):

3x2 - 11x -4 = 0 3x2 -12x +x -4= 0 3x(x-4) +1(x-4) = 0 (3x+1)(x-4) = 0 (3x+1) = 0 , (x-4) = 0 x = -1/3 , x = 4 solution is (-1/3 , 4)